OpenStudy (anonymous):

@Abhisar Please help me... 1)A trolley P of mass m travelling at a speed of 6u towards another trolley Q of mass 4m which is at rest.The trolleys collide elastically.What is the speed of trolley Q after collision?

OpenStudy (anonymous):

2)An alpha particle travelling at a velocity u makes an oblique elastic collision with a helium atom moving with a velocity which negligible compared to u.After collision , the alpha particle moves in the direction inclined at 30 degree to its initial direction of motion.What is the angle between the directions of motion of the helium atom and the initial direction of motion of alpha particle?

OpenStudy (anonymous):

3) A particle of mass m travelling with a velocity u makes a head-on elastic collision with another particle of mass M which is at rest.If m<M,show that the particles with mass m moves in the direction opposite to its initial velocity after collision.

OpenStudy (abhisar):

#1 use \(\large \sf V_2=\frac{2m_1}{m_1+m_2}u_1\\V_2=\frac{2m}{m+4m}6u\)

OpenStudy (abhisar):

Please post one question per thread !